# M3DLayout: A Multi-Source Dataset of 3D Indoor Layouts and Structured Descriptions for 3D Generation

### v1

@@ Abstract
In text-driven 3D scene generation, object layout serves as a crucial intermediate representation that bridges high-level language instructions with detailed geometric output. It not only provides a structural blueprint for ensuring physical plausibility but also supports semantic controllability and interactive editing. However, the learning capabilities of current 3D indoor layout generation models are constrained by the limited scale, diversity, and annotation quality of existing datasets. To address this, we introduce M3DLayout, a large-scale, multi-source dataset for 3D indoor layout generation. M3DLayout comprises 15,080 layouts and over 258k object instances, integrating three distinct sources: real-world scans, professional CAD designs, and procedurally generated scenes. Each layout is paired with detailed structured text describing global scene summaries, relational placements of large furniture, and fine-grained arrangements of smaller items. This diverse and richly annotated resource enables models to learn complex spatial and semantic patterns across a wide variety of indoor environments. To assess the potential of M3DLayout, we establish a benchmark using a text-conditioned diffusion model. Experimental results demonstrate that our dataset provides a solid foundation for training layout generation models. Its multi-source composition enhances diversity, notably through the Inf3DLayout subset which provides rich small-object information, enabling the generation of more complex and detailed scenes. We hope that M3DLayout can serve as a valuable resource for advancing research in text-driven 3D scene synthesis.
Figure 1: The M3DLayout dataset — A multi-source benchmark for text-to-3D indoor scene generation. Top: An example from our dataset showing a detailed 3D indoor layout with richly annotated bounding boxes and its corresponding structured textual description. Bottom-left: Word cloud visualization demonstrating the diversity of room types, furniture, and objects in the dataset. Bottom-right: Overview of the large-scale collection containing 15,080 diverse 3D layout scenes with various styles.
@@ 1 Introduction
Recent advances in 3D generative modeling have enabled remarkable progress in synthesizing 3D objects and scenes from various modalities, such as text or images. These developments have great potential for downstream applications in areas such as content creation, robotics, and virtual reality Höllein et al. (2023); Yang et al. (2024a; b); Schult et al. (2024). In particular, text-to-3D generation has attracted increasing attention due to its intuitive and flexible interface for controlling complex scene content. For example, LucidDreamer Chung et al. (2023) and Text2Immersion Ouyang et al. (2023) adopt point-based or Gaussian-splatting representations to generate detailed scene geometry directly from text or other modalities. Other approaches, such as ATISS Paschalidou et al. (2021), SceneFormer Wang et al. (2021), EchoScene Zhai et al. (2025), and MIDI Huang et al. (2024), perform joint layout-object generation by autoregressively predicting room structures and furnishing objects in a unified framework. LayoutGPT Feng et al. (2023), HoloDeck Yang et al. (2024b), and InstructScene Lin & Mu (2024) employ large language models to plan scene layouts from free-form descriptions, showcasing a promising direction in LLM-driven compositional layout generation.
@@
While these methods demonstrate strong capabilities, they also reveal key limitations. Some of them generate scenes as inseparable volumetric representations, which limits modularity and downstream controllability. Layout-aware models such as CommonScenes Zhai et al. (2023) and DiffuScene Tang et al. (2024) tend to produce relatively simple scenes with few object types and limited diversity. Meanwhile, LLM-based planners show promise in parsing natural language but often struggle with spatial consistency and accurate physical arrangements.
@@
As an essential and intermediate representation of 3D scene, 3D layout data is crucial for high-quality generation. Its importance can be summarized in three key points: $(i)$ 3D layout data define the position, orientation, and scale of objects within a scene, forming the foundation of its structure and spatial coherence. This ensures objects are placed logically and functionally, greatly enhancing the generated scene’s realism and credibility while preventing chaotic or illogical visual outcomes. $(ii)$ As powerful conditional information, 3D layout data guides and constrains the 3D scene generation process. It significantly reduces the model’s degrees of freedom and ambiguity, allowing for more efficient and accurate detail filling. This also enables users to precisely control the scene’s layout, meeting personalized and diverse generation demands. $(iii)$ Real-world 3D scenes typically serve specific functions, and 3D layout data directly reflect this functionality. Through logical arrangements, generated scenes can better fulfill practical application needs (e.g., interior design, game levels) and provide interactive spaces that align with human cognitive habits, thereby significantly improving the end-user experience.
@@
We argue that a major bottleneck limiting further progress in controllable and high-quality scene generation is the lack of large-scale, richly annotated 3D layout datasets that provide structured, semantic-level supervision. Existing datasets either focus on scene geometry from real-world scans (e.g., ScanNet Dai et al. (2017), Matterport3D Chang et al. (2017)) or offer object-level annotations based on professional designs (e.g., 3D-Front Fu et al. (2021), Structured3D Zheng et al. (2020)). However, none of them provide comprehensive layout annotations that include both large-scale furniture and small functional or decorative objects, paired with structured descriptions capturing global scene organization and fine-grained spatial relations.
@@
To address this gap, we introduce M3DLayout, a multi-source dataset for 3D indoor layout generation from structured language. M3DLayout contains 15,080 layouts and over 258k object instances, collected from three complementary sources: real-world scans, professional CAD designs, and procedurally generated scenes. Each layout is paired with structured text descriptions that include global scene summaries, relational placements of large furniture, and fine-grained arrangements of smaller items. These annotations are constructed using a combination of rule-based extraction, GPT-assisted generation, and human verification.
To demonstrate the utility of M3DLayout, we train a text-conditioned diffusion model as a baseline for layout generation. Our experiments show that progressively incorporating the three data sources during training leads to consistent improvements in generation quality, supporting the complementary value of each data type. Furthermore, our model outperforms recent layout generation methods in terms of semantic alignment, spatial plausibility, and controllability.
@@
  1. 1.
    We propose M3DLayout, a large-scale, richly annotated dataset of 3D indoor layouts with structured text descriptions, collected from multiple complementary sources.
  2. 2.
@@
  3. 3.
    We establish a benchmark for text-to-layout generation using a diffusion-based baseline. Results validate the utility of M3DLayout, showing that the dataset enhances the diversity and detail of generated scenes.
@@ 2.1 Datasets for Indoor Scenes
Large-scale datasets play a crucial role in learning-based 3D indoor layout generation and scene synthesis. Early efforts to capture real-world environments using 3D scans led to the creation of datasets such as ScanNet Dai et al. (2017), Matterport3D Chang et al. (2017), and SceneNN Hua et al. (2016). These datasets provide high-fidelity mesh reconstructions, reflect real-world object distributions and spatial constraints. However, they often suffer from noisy geometry, incomplete object coverage, and a lack of fine-grained annotations suitable for generative tasks, as well as limited layout variability due to constrained capture environments.
To address these limitations, synthetic datasets such as SUNCG Song et al. (2017), 3D-FRONT Fu et al. (2021), and Structured3D Zheng et al. (2020) were introduced, offering structured 3D layouts with complete object metadata and annotations from professional CAD designs. However, these professional designs typically lack object variety and fine-grained detail.
Recent hybrid datasets attempt to bridge this gap. FurniScene Zhang et al. (2024) enriches layout realism with more diverse furniture arrangements. OpenRooms Li et al. (2021) provides photorealistic rendering with physical material properties and lighting. Despite these advances, a key limitation persists across nearly all prior datasets: the lack of scene-level textual annotations, which limits their use for conditional or multimodal generation tasks.
@@ 2.2 Indoor Scene Synthesis
The evolution of indoor scene synthesis methods underscores the critical need for richer, more descriptive layout data. Procedural approaches generate indoor scenes using predefined rules, templates, or simulation engines. Systems such as ProcTHOR Deitke et al. (2022) and Infinigen Raistrick et al. (2023; 2024) rely on large-scale procedural scene grammars and asset libraries to create diverse, physically realistic environments. These methods offer high controllability and scalability, especially for generating synthetic training data for embodied agents. However, they are ultimately constrained by their handcrafted rules.
In parallel, a large body of work uses learning-based generative models trained on indoor scene datasets to learn object layouts and spatial relationships in a data-driven fashion. Early methods relied on auto-encoding architectures (e.g., SG-VAE Purkait et al. (2020), SceneHGN Gao et al. (2023), CommonScenes Zhai et al. (2023)) and autoregressive models (e.g., ATISS Paschalidou et al. (2021), SceneFormer Wang et al. (2021)). More recently, diffusion-based models such as DiffuScene Tang et al. (2024) and EchoScene Zhai et al. (2025) have shown superior performance in capturing complex spatial dependencies. SemlayoutDiff Sun et al. (2025)optimized the Diffuscene Tang et al. (2024) by using semantic-mediated 2D distribution maps to finally generate 3D layouts, and is capable of generating different room types with the same model. However, precise control of small object generation cannot be achieved through functional zoning alone. These methods improve the plausibility and diversity of generated layouts but often face challenges in optimization, attribute disentanglement, and generalization to out-of-distribution scenes.
Inspired by the success of large language models (LLMs), several works explore text-guided 3D scene generation. Architect Wang et al. (2024) and HoloDeck Yang et al. (2024b) leverage world knowledge to interpret user instructions and generate spatial constraints. Similarly, FlairGPT Littlefair et al. (2025) controls layout generation through more detailed object descriptions using LLMs, but the overly complex conversational process greatly reduces the efficiency of generating large amounts of data. While these methods support more flexible and intuitive interactions, they often struggle with precise 3D spatial reasoning, often producing physically implausible layouts. This core limitation stems from their lack of grounding in large-scale, physically plausible 3D scene data. M3DLayout is designed to address this critical gap, providing the missing link and a robust foundation to train and benchmark the next generation of these powerful synthesis models.
@@
Figure 3: Dataset statistics of M3DLayout. (a) Top 15 most frequent object categories. (b) Distribution of the number of objects per scene. (c) Proportion of scenes contributed by each source dataset.
@@
Inf3DLayout 7,642 197,707 25.9 57,125 140,582 71.1%
15,080 258,413 17.1 103,391 149,022 57.7%
Table 2: Comparisons between existing 3D indoor scene datasets, where “N/A” denotes “not available”, “L” and “S” denote Large and small objects in the scene, respectively.
@@
SUN3D Xiao et al. (2013) 254 N/A Real scan Low N/A
SceneNNHua et al. (2016) 100 N/A Real scan Low N/A
Matterport3DChang et al. (2017) 1,684 N/A Real scan Medium L-S
ScanNet Dai et al. (2017) 1,506 N/A Real scan Low L-S
Scan2CAD Avetisyan et al. (2019) 1,506 N/A Real scan Low N/A
OpenRooms Li et al. (2021) 1,068 97,607 Real scan Low N/A
SceneNet Handa et al. (2016) 57 3,699 Professional Low N/A
Structured3D Zheng et al. (2020) N/A N/A Professional Low N/A
3D-FRONT Fu et al. (2021) 5,754 N/A Professional Low L
M3DLayout 15,080 258,413 Mixture High L-S
3 The M3DLayout Dataset
To advance controllable and generalizable text-to-3D scene generation, we introduce M3DLayout, a large-scale, multi-source dataset of 3D indoor layouts. This dataset integrates 15,080 layouts from three complementary types of sources: real-world scans, professional interior designs, and procedurally generated scenes. Each layout is annotated with detailed structured textual descriptions to support fine-grained, text-conditioned layout generation.
3.1 Data Sources and Curation
M3DLayout is built upon three distinct types of data sources, each contributing unique characteristics to the dataset:
Real-world Scans. We incorporate layouts from the Matterport3D dataset Chang et al. (2017), which are derived from real environment scans. These layouts reflect realistic, and often cluttered, spatial arrangements found in actual indoor settings. To ensure data quality, we performed a cleanup of the object category list by merging and removing low-frequency categories. Scenes containing fewer than two object instances were filtered out. This curated subset provides a wide variety of scene types and offers important cues for learning robust and realistic layout patterns.
Professional Interior Designs. We integrate high-quality layouts from 3D-FRONT Fu et al. (2021), which contains professionally designed indoor scenes. These layouts are characterized by well-organized spatial semantics and adhere to tidy, minimalist design principles. They typically feature sparser object arrangements, providing strong supervision for generating structurally coherent and aesthetically plausible scenes with clean layouts. Following established practices in prior work Paschalidou et al. (2021); Tang et al. (2024), we applied filters to remove layouts with uncommon object configurations or unnatural room proportions, ensuring a focus on typical, well-structured interior designs.
Procedurally Generated Scenes. To systematically enhance object diversity, particularly for small and decorative items, we generate the Inf3DLayout subset using the procedural generator Infinigen Raistrick et al. (2024). A key part of our curation involved carefully configuring the generator to produce plausible layouts for five common room types: bedrooms, living rooms, dining rooms, kitchens, and bathrooms. The generated houses were then programmatically partitioned into individual rooms. Finally, we applied a filtering step to remove rooms with abnormal layouts or spatial inconsistencies. This curated subset significantly increases the variety and granularity of object arrangements, covering numerous long-tail scenarios that are underrepresented in scan-based or design-based data.
3.2 Structured Description Annotation
To support fine-grained text-conditioned layout generation, we annotate each 3D layout in the M3DLayout dataset with a comprehensive structured description. The annotation schema is designed to capture spatial and semantic information at multiple levels of detail, comprising three key components.

### v2

@@ Abstract
In text-driven 3D scene generation, object layout serves as a crucial intermediate representation that bridges high-level language instructions with detailed geometric output. It not only provides a structural blueprint for ensuring physical plausibility but also supports semantic controllability and interactive editing. However, the learning capabilities of current 3D indoor layout generation models are constrained by the limited scale, diversity, and annotation quality of existing datasets. To address this, we introduce M3DLayout, a large-scale, multi-source dataset for 3D indoor layout generation. M3DLayout comprises 21,367 layouts and over 433k object instances, integrating three distinct sources: real-world scans, professional CAD designs, and procedurally generated scenes. Each layout is paired with detailed structured text describing global scene summaries, relational placements of large furniture, and fine-grained arrangements of smaller items. This diverse and richly annotated resource enables models to learn complex spatial and semantic patterns across a wide variety of indoor environments. To assess the potential of M3DLayout, we establish a benchmark using both a text-conditioned diffusion model and a text-conditioned autoregressive model. Experimental results demonstrate that our dataset provides a solid foundation for training layout generation models. Its multi-source composition enhances diversity, notably through the Inf3DLayout subset which provides rich small-object information, enabling the generation of more complex and detailed scenes. We hope that M3DLayout can serve as a valuable resource for advancing research in text-driven 3D scene synthesis. All dataset and code will be made public upon acceptance.
@@ 1 Introduction
Recent advances in 3D generative modeling have enabled remarkable progress in synthesizing 3D objects and scenes from various modalities, such as text or images. These developments have great potential for downstream applications in areas such as content creation, robotics, and virtual reality [holleinText2RoomExtractingTextured2023, yangDreamSpaceDreamingYour2024, yangHolodeckLanguageGuided2024a, schultControlRoom3DRoomGeneration2024]. In particular, text-to-3D generation has attracted increasing attention due to its intuitive and flexible interface for controlling complex scene content. For example, LucidDreamer [chungLucidDreamerDomainfreeGeneration2023] and Text2Immersion [ouyangText2ImmersionGenerativeImmersive2023] adopt point-based or Gaussian-splatting representations to generate detailed scene geometry directly from text or other modalities. Other approaches, such as ATISS [paschalidouATISSAutoregressiveTransformers2021], SceneFormer [wangSceneFormerIndoorScene2021], EchoScene [zhaiEchoSceneIndoorScene2025], and MIDI [huangMIDIMultiInstanceDiffusion2024], perform joint layout-object generation by autoregressively predicting room structures and furnishing objects in a unified framework. LayoutGPT [fengLayoutGPTCompositionalVisual2023], HoloDeck [yangHolodeckLanguageGuided2024a], and InstructScene [linInstructSceneInstructionDriven3D2024] employ LLMs to plan scene layouts from free-form descriptions, showcasing a promising direction in LLM-driven compositional layout generation.
@@
While these methods demonstrate strong capabilities, they also reveal key limitations. Some of them generate scenes as inseparable volumetric representations, which limits modularity and downstream controllability. Layout-aware models such as CommonScenes [zhaiCommonScenesGeneratingCommonsense2023] and DiffuScene [tangDiffuSceneDenoisingDiffusion2024] tend to produce relatively simple scenes with few object types and limited diversity. Meanwhile, LLM-based planners show promise in parsing natural language but often struggle with spatial consistency and accurate physical arrangements.
To overcome this bottleneck, we draw inspiration from AutoPartGen [chen2025autopartgen], which decomposes a scene into distinct parts and thus establishes a bridge between part-level object generation and scene generation. However, AutoPartGen is restricted to small and relatively simple scenes since they simply operate under masked-image conditioning, making it difficult to scale to complex 3D scene environments. In contrast, mainstream object-level part generation methods such as OmniPart [yang2025omnipart], X-Part [yan2025x], and CoPart [dong2025copart] rely heavily on 3D bounding boxes as context to control the locations and numbers of parts, highlighting the crucial role of 3D bounding boxes as an intermediate representation. When these 3D bounding boxes are further enriched with semantic labels, rotation information and text descriptions, they form a full 3D layout that not only defines the structural backbone of the scene, but also provides powerful conditional guidance and reflects the functional intent of the environment. Building on this insight, we integrate such layout representations into the scene generation pipeline to produce high-quality, controllable 3D scenes with coherent spatial structure and realistic functionality and further unify object generation and scene generation.
@@
To be more specific, the importance of 3D layouts can be summarized in three key points: $(i)$ 3D layout data define the position, orientation, and scale of objects within a scene, forming the foundation of its structure and spatial coherence. This ensures objects are placed logically and functionally, greatly enhancing the generated scene’s realism and credibility while preventing chaotic or illogical visual outcomes. $(ii)$ As powerful conditional information, 3D layout data guides and constrains the 3D scene generation process. It significantly reduces the model’s degrees of freedom and ambiguity, allowing for more efficient and accurate detail filling. This also enables users to precisely control the scene’s layout, meeting personalized and diverse generation demands. $(iii)$ Real-world 3D scenes typically serve specific functions, and 3D layout data directly reflect this functionality. Through logical arrangements, generated scenes can better fulfill practical application needs (e.g., interior design, game levels) and provide interactive spaces that align with human habits, thereby significantly improving the end-user experience.
@@
We argue that a major constraint limiting further progress in controllable and high-quality scene generation is the lack of large-scale, richly annotated 3D layout datasets that provide structured, semantic-level supervision. Existing datasets either focus on scene geometry from real-world scans (e.g., ScanNet [daiScanNetRichlyAnnotated3D2017a], Matterport3D [changMatterport3DLearningRGBD2017]) or offer object-level annotations based on professional designs (e.g., 3D-FRONT [fu3DFRONT3DFurnished2021], Structured3D [zhengStructured3DLargePhotorealistic2020]). However, none of them offer comprehensive layout annotations—covering large furniture, small objects, and decorative items—paired with structured descriptions of global scene organization and fine-grained spatial relations, and scene feasibility issues, such as overlapping, displacement, and semantic violations, are rarely verified, leaving substantial noise that ultimately limits text-to-layout generation and scene generation.
@@
To address this gap, we introduce M3DLayout, a multi-source dataset for 3D indoor layout generation from structured language. M3DLayout contains 21,367 layouts and over 433k object instances, collected from three complementary sources: real-world scans, professional CAD designs, and procedurally generated scenes. We pair each layout with structured text descriptions that capture global scene organization, relational placements of large furniture, and fine-grained arrangements of smaller items. These high-quality annotations were constructed using a combination of rule-based extraction, GPT-assisted generation, and rigorous human verification.
To demonstrate the utility of M3DLayout, we establish a benchmark using both a text-conditioned diffusion model and a text-conditioned autoregressive model as baselines for layout generation. Experimental results show that our dataset provides a solid foundation for training layout generation models. Its multi-source composition enhances diversity, notably through the Inf3DLayout subset which provides rich small-object information, enabling the generation of more complex and detailed scenes.
@@
  • We introduce M3DLayout, a large-scale, richly annotated dataset of 3D indoor layouts with structured text descriptions, compiled from diverse complementary sources.
@@
  • We establish two baselines for text-to-layout generation using both diffusion-based and autoregressive baselines. Results show that M3DLayout enhances the diversity and detail of generated scenes, and also enables explicit control over the fineness and density of the generated 3D layout via textual prompts.
@@ 2.1 Datasets for Indoor Scenes
Large-scale datasets play a crucial role in learning-based 3D indoor layout generation and scene synthesis. Early efforts to capture real-world environments using 3D scans led to the creation of datasets such as ScanNet [daiScanNetRichlyAnnotated3D2017a], Matterport3D [changMatterport3DLearningRGBD2017], and SceneNN [huaSceneNNSceneMeshes2016]. These datasets provide high-fidelity mesh reconstructions, reflect real-world object distributions and spatial constraints. However, they often suffer from noisy geometry, incomplete object coverage, and a lack of fine-grained annotations suitable for generative tasks, as well as limited layout variability due to constrained capture environments.
To address these limitations, synthetic datasets such as SUNCG [songSemanticSceneCompletion2017], 3D-FRONT [fu3DFRONT3DFurnished2021], and Structured3D [zhengStructured3DLargePhotorealistic2020] were introduced, offering structured 3D layouts with complete object metadata and annotations from professional CAD designs. However, these professional designs typically lack object variety and fine-grained detail.
Recent hybrid datasets attempt to bridge this gap. FurniScene [zhangFurniSceneLargescale3D2024] enriches layout realism with more diverse furniture arrangements. OpenRooms [liOpenRoomsEndtoEndOpen2021] provides photorealistic rendering with physical material properties. However, a key limitation persists across nearly all prior datasets: the lack of scene-level textual annotations, which limits their use for conditional or multimodal generation tasks.
@@ 2.2 Indoor Scene Synthesis
The evolution of indoor scene synthesis methods highlights the need for richer, more descriptive layout data. Procedural approaches generate indoor scenes using predefined rules, templates, or simulation engines. Systems such as ProcTHOR [deitkeProcTHORLargeScaleEmbodied2022] and Infinigen [raistrickInfinitePhotorealisticWorlds2023a, raistrickInfinigenIndoorsPhotorealistic2024] rely on large-scale procedural scene grammars and asset libraries to create diverse, physically realistic environments. These methods offer high controllability and scalability, especially for generating synthetic training data for embodied agents. However, they are ultimately constrained by their handcrafted rules.
In parallel, a large body of work uses learning-based generative models trained on indoor scene datasets to learn object layouts and spatial relationships in a data-driven fashion. Early methods relied on auto-encoding architectures (e.g., SG-VAE [purkaitSGVAESceneGrammar2020], SceneHGN [gaoSceneHGNHierarchicalGraph2023], CommonScenes [zhaiCommonScenesGeneratingCommonsense2023]) and autoregressive models (e.g., ATISS [paschalidouATISSAutoregressiveTransformers2021], SceneFormer [wangSceneFormerIndoorScene2021]). More recently, diffusion-based models such as DiffuScene [tangDiffuSceneDenoisingDiffusion2024] and EchoScene [zhaiEchoSceneIndoorScene2025] have shown superior performance in capturing complex spatial dependencies. SemlayoutDiff [sun_semlayoutdiff_2025]optimized the Diffuscene [tangDiffuSceneDenoisingDiffusion2024] by using semantic-mediated 2D distribution maps to finally generate 3D layouts, and is capable of generating different room types with the same model. However, precise control of small object generation cannot be achieved through functional zoning alone. These methods improve the plausibility and diversity of generated layouts but often face challenges in optimization, attribute disentanglement, and generalization to out-of-distribution scenes.
Inspired by the success of large language models (LLMs), several works explore text-guided 3D scene generation. Architect [wangArchitectGeneratingVivid2024] and HoloDeck [yangHolodeckLanguageGuided2024a] leverage world knowledge to interpret user instructions and generate spatial constraints. Similarly, FlairGPT [littlefair_flairgpt_2025] controls layout generation through more detailed object descriptions using LLMs, but the complex conversational process greatly reduces the efficiency of generating large amounts of data. While these methods support flexible and intuitive interactions, they often struggle with precise 3D spatial reasoning, often producing physically implausible layouts. This core limitation stems from their lack of grounding in large-scale, physically plausible 3D scene data. M3DLayout is designed to address this critical gap, providing the missing link and a robust foundation to train and benchmark the next generation of these powerful synthesis models.
@@
Figure 3: Dataset statistics of M3DLayout. (a) Top 15 most frequent object categories. (b) Distribution of the number of objects per scene. (c) Proportion of scenes contributed by each source.
3 The M3DLayout Dataset
To advance controllable and generalizable text-to-3D scene generation, we introduce M3DLayout, a large-scale, multi-source dataset of 3D indoor layouts. This dataset integrates 21,367 layouts from three complementary types of sources: real-world scans, professional interior designs, and procedurally generated scenes. Each layout is annotated with detailed structured textual descriptions to support fine-grained, text-conditioned layout generation.
3.1 Data Sources and Curation
M3DLayout is built upon three distinct types of data sources, each contributing unique characteristics:
Real-world Scans. We incorporate layouts from the Matterport3D dataset [changMatterport3DLearningRGBD2017], which are derived from real environment scans. These layouts reflect realistic, and often cluttered, spatial arrangements found in actual indoor settings. To ensure data quality, we performed a cleanup of the object category list by merging and removing low-frequency categories. Scenes containing fewer than two object instances were filtered out. This curated subset provides a wide variety of scene types and offers important cues for learning robust and realistic layout patterns.
Professional Interior Designs. We integrate high-quality layouts from 3D-FRONT [fu3DFRONT3DFurnished2021], which contains professionally designed indoor scenes. These layouts are characterized by well-organized spatial semantics and adhere to tidy, minimalist design principles. They typically feature sparser object arrangements, providing strong supervision for generating structurally coherent and aesthetically plausible scenes with clean layouts. Following prior works [paschalidouATISSAutoregressiveTransformers2021, tangDiffuSceneDenoisingDiffusion2024], we applied filters to remove layouts with uncommon object configurations or unnatural room proportions, ensuring a focus on typical and well-structured designs.
Procedurally Generated Scenes. To systematically enhance object diversity, particularly for small and decorative items, we generate the Inf3DLayout subset using the procedural generator Infinigen [raistrickInfinigenIndoorsPhotorealistic2024]. A key part of our curation involved carefully configuring the generator to produce plausible layouts for five common room types: bedrooms, living rooms, dining rooms, kitchens, and bathrooms. The generated houses were then programmatically partitioned into individual rooms. Finally, we applied a filtering step to remove rooms with abnormal layouts or spatial inconsistencies. This curated subset significantly increases the variety and granularity of object arrangements, covering numerous long-tail scenarios that are underrepresented in scan-based or design-based data.
3.2 Structured Description Annotation
To support fine-grained text-conditioned layout generation, we annotate each 3D layout in the M3DLayout dataset with a comprehensive structured description. The annotation schema is designed to capture spatial and semantic information at multiple levels, comprising 3 key components.
@@ 3.2 Structured Description Annotation
Inf3DLayout 13,929 373,136 26.8 117,700 255,426 68.5%
21,367 433,842 20.3 169,966 263,866 60.8%
Table 2: Comparisons between existing 3D indoor scene datasets. For the column of Layout Collecton, RS denotes real scanned and PD denotes professionally designed .For the column of Variation in Object Sizes, “N/A” denotes “not available”, “L” and “S” denote Large and Small objects in the scene, respectively.
@@
SUN3D [xiao2013sun3d] 254 N/A RS Low N/A
SceneNN[huaSceneNNSceneMeshes2016] 100 N/A RS Low N/A
Matterport3D[changMatterport3DLearningRGBD2017] 1,684 N/A RS Medium L-S
ScanNet [daiScanNetRichlyAnnotated3D2017a] 1,506 N/A RS Low L-S
Scan2CAD [avetisyan2019scan2cad] 1,506 N/A RS Low N/A
OpenRooms [liOpenRoomsEndtoEndOpen2021] 1,068 97,607 RS Low N/A
SceneNet [handa2016understanding] 57 3,699 PD Low N/A
Structured3D [zhengStructured3DLargePhotorealistic2020] N/A N/A PD Low N/A
3D-FRONT [fu3DFRONT3DFurnished2021] 5,754 N/A PD Low L
M3DLayout (Ours) 21,367 433,842 Mixture High L-S